# How does the successive buckling events affect a galaxy bar and stellar disk? : Potential Observable Signatures For Spotting the Buckling Action -I

### v1

@@ 5 Conclusions
  • First buckling shows a glitch in the time evolution of the pattern speed while second buckling does not show.
  • First buckling with a short time span leads to a shorter BPX bulge ($\approx 5.8$ kpc) and the second buckling with a larger time span leads to a longer BPX bulge ($\approx$ 8.6 kpc). BPX bulge lengths corresponding to successive buckling events are around two-thirds of the bar lengths at the end of individual buckling events.
  • The kinematic signature for the bar buckling is observed in the Los mean velocity for a nearly face-on galaxy disk. Both the buckling shows the quadrupolar patterns in mean Los velocity which last thoughout the buckling event. The quadrupolar patterns in the first buckling are compact, symmetric, and pronounced compared to those in the second buckling, which are extended, non-symmetric, and less pronounced.
@@
  • First buckling leads to a sharp dip in radial dispersion along the bar and a sharp rise in vertical dispersion throughout the disk, suggesting instant transfer of energy from a horizontal plane to the vertical plane of the disk. The second buckling leads to the flattening of radial dispersion along the bar and a slow increase in vertical velocity dispersion throughout the disk, suggesting a relatively slower transfer of energy across horizontal and vertical planes.
@@ Acknowledgements
The research presented here is partially supported by the National Key R&D Program of China under grant No. 2018YFA0404501; by the National Natural Science Foundation of China under grant Nos. 12025302, 11773052, 11761131016; by the “111” Project of the Ministry of Education of China under grant No. B20019; and by the China Manned Space Project under grant No. CMS-CSST-2021-B03. SKK thanks Volker Springel for the Gadget code which we used to run our simulations. SKK thanks Juntai Shen, Zhi Li, Rui Guo, Xingchen Li and Yirui Zheng for the scientific discussion which helped to progress in this work. This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University, and the facilities of the Center for High Performance Computing at Shanghai Astronomical Observatory. software used: numpy (Harris et al., 2020), matplotlib (Hunter, 2007), pynbody (Pontzen et al., 2013) and astropy (Astropy Collaboration et al., 2018) packages.

### v2

@@ 5 Conclusions
  • First buckling shows an abrupt change in the time evolution of the pattern speed while second buckling does not show this. This abrupt change in pattern speed leads to a faster decrease in pattern speed followed by sudden increase during the buckling event.
  • First buckling with a short time span leads to a shorter BPX bulge ($\approx 5.8$ kpc) and the second buckling with a larger time span leads to a longer BPX bulge ($\approx$ 8.6 kpc). BPX bulge full lengths corresponding to successive buckling events are around two-thirds of the full bar lengths at the end of individual buckling events.
  • The kinematic signature for the bar buckling is observed in the LOS mean velocity for a nearly face-on galaxy disk. Both the buckling events show the quadrupolar patterns in mean LOS velocity which last thoughout the buckling event. The quadrupolar patterns in the first buckling are compact, symmetric, and pronounced compared to those in the second buckling, which are extended, non-symmetric, and less pronounced.
@@
  • First buckling leads to a sharp dip in radial dispersion along the bar and a sharp rise in vertical dispersion throughout the disk, suggesting very rapid transfer of energy from a horizontal plane to the vertical plane of the disk. The second buckling leads to the flattening of radial dispersion along the bar and a slow increase in vertical velocity dispersion throughout the disk, suggesting a relatively slower transfer of energy across horizontal and vertical planes.
@@ Acknowledgements
SKK thanks anonymous referee for the insightful comments to improve the content in the draft. The research presented here is partially supported by the National Key R&D Program of China under grant No. 2018YFA0404501; by the National Natural Science Foundation of China under grant Nos. 12025302, 11773052, 11761131016; by the “111” Project of the Ministry of Education of China under grant No. B20019; and by the China Manned Space Project under grant No. CMS-CSST-2021-B03. SKK thanks Volker Springel for the Gadget code which we used to run our simulations. SKK thanks Juntai Shen, Zhi Li, Rui Guo, Xingchen Li and Yirui Zheng for the scientific discussion which helped to progress in this work. This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University, and the facilities of the Center for High Performance Computing at Shanghai Astronomical Observatory. software used: numpy (Harris et al., 2020), matplotlib (Hunter, 2007), pynbody (Pontzen et al., 2013) and astropy (Astropy Collaboration et al., 2018) packages.